OpenStudy (anonymous):

SAT question Test 10 Section 8 No. 11: 4(2^X)=2^y. What is x in terms of Y? Can anyone help? thanks!!

OpenStudy (anonymous):

\[4(2^x)=2^2(2^x)=2^{x+2}\]

OpenStudy (anonymous):

so we now have:\[2^{x+2}=2^y\Longrightarrow x+2=y\iff x=y-2\]

OpenStudy (anonymous):

Thanks so much -- i got it. I was first thinking to multiply the 4 and the 2 to make 8^x, then say 2^3+x, then say 3+x=Y. But thats wrong. Not sure why it doesn't work that way though....thanks again!